# An Operator Splitting Method for Large-Scale CVaR-Constrained Quadratic Programs

### v3

@@ Results.
Figure 1 shows solve times versus number of scenarios $m$ for $\tau=0.5$, averaged over 10 random instances. Our specialized algorithm is at least 100x faster than both general-purpose solvers across all problem sizes, with times ranging from $5.08\times 10^{-4}$ seconds at $m=10^{4}$ to $1.98$ seconds at $m=10^{7}$. We achieve up to 828× speedup over Mosek before it fails on larger problems ($m\geq 3\times 10^{6}$). Compared to Clarabel, our method is at least 242x faster across all problem sizes, with maximum speedup of 654x.
Figure 1: CVaR projection solve times versus number of scenarios $m$, with $\tau=0.5$.
@@ 5.3 Portfolio optimization
In portfolio optimization, we allocate capital across $n$ assets to maximize risk-adjusted returns. The decision variable $x\in{\mbox{\bf R}}^{n}$ represents portfolio weights.
We model uncertainty using $m$ return scenarios. The matrix $R\in{\mbox{\bf R}}^{m\times n}$ contains asset returns, where $R_{ij}$ is the return of asset $j$ in scenario $i$. The mean return vector $\mu\in{\mbox{\bf R}}^{n}$ and covariance matrix $\Sigma$ are given by
@@
where $\gamma>0$ is the risk aversion parameter. The first constraint normalizes the portfolio weights, the second prevents short positions, and the CVaR constraint bounds the expected loss in the worst $(1-\beta)$ fraction of scenarios.
@@
$R_{i}\sim\alpha\mathcal{N}(\nu,I)+(1-\alpha)\mathcal{N}(-\nu,\sigma^{2}I),\quad i=1,\ldots,m,$
where $\alpha$ is the probability of normal market conditions, $\nu$ is the mean return vector (positive in normal conditions, negative in stress periods), and $\sigma>1$ scales the volatility during stress periods. We use parameters
$\alpha$ $=0.8,$ $\nu$ $=0.2,$ $\gamma$ $=1,$
$\sigma$ $=2,$ $\beta$ $=0.95,$ $\kappa$ $=0.3.$
@@ Results.
Figure 2 shows solve times versus number of scenarios $m$ for problems with $n=2000$ assets, averaged over 3 random instances. Our method is substantially faster than both general-purpose solvers. At $m=10^{4}$ scenarios, we are 140× faster than Mosek, maintaining at least a 30× speedup until Mosek fails at $m=10^{6}$. Our advantage over Clarabel is even larger, reaching 2,914× at $m=3\times 10^{4}$ scenarios before Clarabel fails at $m\geq 10^{5}$. Our method solves problems with up to one million scenarios in under 26 minutes, when others fail.
Figure 2: Solve times versus number of scenarios $m$ for portfolio optimization problem.
@@ 5.4 Quantile regression
In quantile regression, we estimate the $\tau$-quantile of a response variable conditional on features. Given $m$ data points with features $u_{i}\in{\mbox{\bf R}}^{n}$ and responses $y_{i}\in{\mbox{\bf R}}$, the problem is
$\begin{array}[]{ll}\text{minimize}&\frac{1}{m}\sum_{i=1}^{m}\rho_{\tau}(y_{i}-x^{T}u_{i}-x_{0}),\end{array}$
with variables $x\in{\mbox{\bf R}}^{n}$ and $x_{0}\in{\mbox{\bf R}}$. Here, $\rho_{\tau}(z)=\tau(z)_{+}+(1-\tau)(z)_{-}$ is the tilted $\ell_{1}$ penalty function, where $(z)_{+}=\max\{z,0\}$ and $(z)_{-}=\max\{-z,0\}$. The parameter $\tau\in(0,1)$ controls the quantile level, with $\tau=0.5$ giving median regression.
@@
We can express the tilted $\ell_{1}$ penalty as $\rho_{\tau}(z)=(1-\tau)(-z)+(z)_{+}$, which after scaling by $(1-\tau)^{-1}$ gives the objective
$x_{0}+\frac{1}{m(1-\tau)}\sum_{i=1}^{m}\max\{y_{i}-x^{T}u_{i}-x_{0},0\}-\frac{1}{m}\sum_{i=1}^{m}(y_{i}-x^{T}u_{i}).$
The first term is precisely the CVaR at level $\tau$ of the residuals. With feature matrix $U\in{\mbox{\bf R}}^{m\times n}$ with rows $u_{i}^{T}$ and response vector $y\in{\mbox{\bf R}}^{m}$, we get the equivalent problem
$\begin{array}[]{ll}\text{minimize}&x^{T}\bar{u}+t\\ \text{subject to}&\phi_{\tau}(-Ux+y)\leq t,\end{array}$
where $\bar{u}=\frac{1}{m}\sum_{i=1}^{m}u_{i}$ and $t$ is an auxiliary variable.
@@
To express this in our standard CVQP form, we introduce an auxiliary variable $s$ with the constraint $s=1$, giving
$\begin{array}[]{ll}\text{minimize}&x^{T}\bar{u}+t\\ \text{subject to}&\phi_{\tau}(-Ux-t\mathbf{1}+sy)\leq 0\\ &s=1.\end{array}$
Defining the variable $\tilde{x}=(x,t,s)\in{\mbox{\bf R}}^{n+2}$, we can express this in our standard CVQP form with parameters
$P$ $=0,$ $q$ $=\begin{bmatrix}\bar{u}\\ 1\\ 0\end{bmatrix},$ $\kappa$ $=0,$
$A$ $=\begin{bmatrix}-U&-\mathbf{1}&y\end{bmatrix},$ $B$ $=\begin{bmatrix}0^{T}&0&1\end{bmatrix},$ $l=u$ $=1$
The matrix $A\in{\mbox{\bf R}}^{m\times(n+2)}$ contains the coefficients of $\tilde{x}$ in the expression $-Ux-t\mathbf{1}+sy$. The constraint matrix $B\in{\mbox{\bf R}}^{1\times(n+2)}$ extracts the $s$ component, ensuring $s=1$.
@@ Results.
Figure 3 shows solve times versus number of scenarios $m$ for quantile regression problems with $n=500$ features. Our method outperforms both general-purpose solvers across all problem sizes. At $m=10^{4}$ scenarios, our method is 17x faster than Mosek and 178x faster than Clarabel. At $m=10^{6}$, we maintain a 4.6x advantage over Mosek before it fails on larger problems. Clarabel cannot handle problems beyond $m=3\times 10^{5}$. Our approach scales to $m=10^{7}$ scenarios with consistent performance
Figure 3: Solve times versus number of scenarios $m$ for quantile regression problem.
@@ 6 Extensions and variations
Our method extends naturally to handle more general problems and admits several variations that improve performance.
@@
ADMM extends naturally to more general constraints on $\tilde{z}$. When projection onto a set $\mathcal{C}$ is efficient (via a closed-form solution or a fast algorithm), we can directly handle constraints of the form $\tilde{z}\in\mathcal{C}$, allowing efficient solution of a broader class of problems.
@@
Our method can be extended to accommodate additional constraints on $x$ or different objective functions. The $x$-update then becomes
@@
Equilibration as a preprocessing step often improves ADMM convergence, particularly for problems with poorly scaled data. While not needed in our benchmark examples, for real-world problems with heterogeneous scales, equilibration can significantly reduce the number of iterations required.
@@
Well-chosen starting points can accelerate convergence. Two effective strategies are: (1) solving with a reduced set of scenarios, then using that solution to initialize the full problem; and (2) using a quadratic approximation of the CVaR constraint to compute an initial point. These approaches can substantially reduce computation time for large-scale problems.

### v4

@@ Results.
Figure 1 shows solve times versus number of scenarios $m$ for $\eta=0.5$, averaged over 10 random instances. Our method solves these problems in $0.51$ ms at $m=10^{4}$ and $1.98$ s at $m=10^{7}$, more than two orders of magnitude faster than both Mosek and Clarabel across all tested sizes. Mosek does not solve instances with $m\geq 3\times 10^{6}$ within the time limit.
Figure 1: CVaR projection solve times, with $\eta=0.5$.
@@ 5.3 Portfolio optimization
We consider a portfolio optimization problem with $n$ assets and $m$ return scenarios. The matrix $R\in{\mbox{\bf R}}^{m\times n}$ contains asset returns, where $R_{ij}$ is the return of asset $j$ in scenario $i$. The mean return vector $\mu\in{\mbox{\bf R}}^{n}$ and covariance matrix $\Sigma$ are given by
@@
with variable $x\in{\mbox{\bf R}}^{n}$ (the portfolio weights), risk aversion parameter $\gamma>0$, and a CVaR constraint bounding the expected loss in the worst $(1-\beta)$ fraction of scenarios.
@@
$R_{i}\sim\omega\mathcal{N}(\nu\mathbf{1},I)+(1-\omega)\mathcal{N}(-\nu\mathbf{1},\sigma^{2}I),\quad i=1,\ldots,m,$
where $\omega$ is the probability of normal market conditions, $\nu$ is the mean return per asset (positive in normal conditions, negative in stress periods), and $\sigma>1$ scales the volatility during stress periods. We use parameters
[…]
@@ Results.
Figure 2 shows solve times versus number of scenarios $m$ for problems with $n=2000$ assets, averaged over 3 random instances. Our method is faster than both general-purpose solvers by one to three orders of magnitude. Mosek does not solve instances with $m\geq 10^{6}$ within the time limit, and Clarabel does not solve instances with $m\geq 10^{5}$. Our method solves instances with up to one million scenarios in under 26 minutes.
Figure 2: Portfolio optimization solve times, with $n=2000$ assets.
@@ 5.4 Quantile regression
Given $m$ data points with features $u_{i}\in{\mbox{\bf R}}^{n}$ and responses $y_{i}\in{\mbox{\bf R}}$, the $\tau$-quantile regression problem is
$\text{minimize}\quad\frac{1}{m}\sum_{i=1}^{m}\rho_{\tau}(y_{i}-x^{T}u_{i}-x_{0}),$
with variables $x\in{\mbox{\bf R}}^{n}$ and $x_{0}\in{\mbox{\bf R}}$, where $\rho_{\tau}(z)=\tau(z)_{+}+(1-\tau)(z)_{-}$ is the tilted $\ell_{1}$ penalty and $\tau\in(0,1)$ is the quantile level. The asymmetric penalty weighs underpredictions $\tau/(1-\tau)$ times more than overpredictions, so the optimal prediction is the conditional $\tau$-quantile of $y$ given $u$, rather than the conditional mean.
@@
Defining the feature matrix $U\in{\mbox{\bf R}}^{m\times n}$ with rows $u_{i}^{T}$, the response vector $y\in{\mbox{\bf R}}^{m}$, and $\bar{u}=\frac{1}{m}\sum_{i=1}^{m}u_{i}$, we use the identity $\rho_{\tau}(z)=(1-\tau)(-z)+(z)_{+}$ to write the objective as
$(1-\tau)\left(x^{T}\bar{u}+x_{0}-\bar{y}\right)+\frac{1}{m}\sum_{i=1}^{m}(y_{i}-u_{i}^{T}x-x_{0})_{+},$
where $\bar{y}=\frac{1}{m}\sum_{i=1}^{m}y_{i}$. The terms involving $x_{0}$ can be written as
$(1-\tau)\left[x_{0}+\frac{1}{(1-\tau)m}\sum_{i=1}^{m}(y_{i}-u_{i}^{T}x-x_{0})_{+}\right],$
which is $(1-\tau)$ times the sample CVaR formula evaluated at $\alpha=x_{0}$. Minimizing over $x_{0}$ yields the infimum in the CVaR definition. Since $\bar{y}$ is constant and $(1-\tau)>0$, the problem reduces to
$\text{minimize}\quad x^{T}\bar{u}+\phi_{\tau}(-Ux+y),$
with variable $x\in{\mbox{\bf R}}^{n}$. Introducing an epigraph variable $t$ for the CVaR term gives
$\begin{array}[]{ll}\text{minimize}&x^{T}\bar{u}+t\\ \text{subject to}&\phi_{\tau}(-Ux+y)\leq t.\end{array}$
@@
We introduce an auxiliary variable $s$ with the constraint $s=1$, and define $\tilde{x}=(x,t,s)\in{\mbox{\bf R}}^{n+2}$. The CVQP parameters, with CVaR level $\beta=\tau$, are
$P$ $=0,$ $q$ $=\begin{bmatrix}\bar{u}\\ 1\\ 0\end{bmatrix}$, $A$ $=\begin{bmatrix}-U&-\mathbf{1}&y\end{bmatrix}$,
$B$ $=\begin{bmatrix}0^{T}&0&1\end{bmatrix}$, $l$ $=1,$ $u$ $=1.$
Problem instances.
We generate features $u_{i}\in{\mbox{\bf R}}^{n}$ with independent standard normal entries, a vector with entries $\beta_{j}\sim\mathcal{N}(0,1/(1+j))$, and responses $y_{i}=u_{i}^{T}\beta+\epsilon_{i}$, where $\epsilon_{i}$ follows a $t$-distribution with 5 degrees of freedom scaled by $0.1$. We use quantile level $\tau=0.9$.
@@ Results.
Figure 3 shows solve times versus number of scenarios $m$ for quantile regression problems with $n=500$ features, averaged over 3 random instances. Our method is faster than both Mosek and Clarabel across all tested sizes, by a factor ranging from about 5 to over 100. Mosek does not solve instances with $m>10^{6}$ within the time limit, and Clarabel does not solve instances with $m\geq 3\times 10^{5}$. Our method scales to $m=10^{7}$ scenarios.
Figure 3: Quantile regression solve times, with $n=500$ features.
@@ 6 Extensions and variations
Our method extends to more general problems and admits several practical variations.
@@
When projection onto a set $\mathcal{C}$ is efficient (via a closed-form solution or a fast algorithm), we can directly handle constraints of the form $\tilde{z}\in\mathcal{C}$ in place of the box constraints $l\leq\tilde{z}\leq u$.
@@
The method extends to additional constraints on $x$ or nonquadratic objective functions. The $x$-update then becomes
@@
Equilibration as a preprocessing step often improves ADMM convergence, particularly for problems with poorly scaled data.
@@
Two effective strategies for choosing the initial point are (1) solving with a reduced set of scenarios and using that solution to initialize the full problem, and (2) using a quadratic approximation of the CVaR constraint to compute an initial point. Similarly, the sorting step in the CVaR projection can be warm-started using the sorted order from the previous ADMM iteration. Since the projection input changes only slightly between iterations, the sorted order is nearly preserved, and an adaptive sorting algorithm can exploit this to reduce the per-iteration cost in practice.